OpenStudy (anonymous):

A train is moving 80ft in 16 seconds, How fast is the train moving in MPH?

OpenStudy (mortonsalt):

v = d/t d = 80 ft = 0.0151515... mi t = 16 seconds = approx 0.004444 h

OpenStudy (anonymous):

Velocity= Distance traveled/ Time taken Distance traveled= 80 ft = (80 x (1/5280)) Miles = 0.01515 Miles 5280 ft = 1 mile Time taken= 16 sec= (16 x (1/(60 x 60))) = 0.00444 Hour 1 hour= 60 min x 60 sec Velocity in MPH = 0.01515/ 0.00444 = 3.4121 MPH

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Its my pleasure! :)